OpenStudy (anonymous):

The area of the trapezoid is 24 and the bases are 3 and 5. Find the height. 3 3.2 6

OpenStudy (anonymous):

the area for a trapeezoid is b1+b2/2 times the height 24=3+5/2 times h 24+8/2 times h 24=4h 6=h

OpenStudy (anonymous):

oh okay thank you i was unsure on how to do the area